### v2

# Reorganizing attention-space geometry with expressive attention

Abstract
Attention regulates information transfer between tokens. For this, query and key vectors are compared, typically in terms of a scalar product, $\mathbf{Q}^{T}\mathbf{K}$, together with a subsequent softmax normalization. In geometric terms, the standard dot-product attention (DPA) leads to large/small attention weights for parallel/antiparallel queries and keys. Here we study expressive attention (EA), which is based on $(\mathbf{Q}^{T}\mathbf{K})^{2}$, the squared dot product. In this case, attention is enhanced when query and key are either parallel or antiparallel, and suppressed for orthogonal configurations. EA can be introduced into any attention-based code without additional compute costs or memory requirements. For a series of autoregressive prediction tasks, we find that expressive attention performs at least as well as vanilla DPA. Increasing task complexity, EA is observed to outperform DPA with increasing margins, which also holds for multi-task settings. For a given model size, EA manages to achieve 100% performance for a range of complexity levels not accessible to DPA. Our results show that it is possible to reorganize the geometry of the matching condition in the space of attention heads without loss of performance.
1 Introduction
Since its inception (?, ?), the attention mechanism has had a transformative impact on machine learning (?, ?). At its core, attention facilitates pairwise information routing between tokens, with information being transmitted when a given matching condition is fulfilled. For the latter, query and key vectors $\mathbf{Q}$ and $\mathbf{K}$ are compared in terms of the respective dot products, $\mathbf{Q}^{T}\mathbf{K}$. In the standard formulation, denoted here dot-product attention (DPA), a softmax operation is used to map the scalar products between queries and keys to a normalized probability distribution. As illustrated in Fig. 1, dot-product attention constrains the matching condition essentially to a one-dimensional subspace of the otherwise potentially large space of attention heads. Given that attention matrices tend to be sparse (?), it may however be favorable to allow the system to express low attention states in a larger subspace.
Here we introduce and discuss a modified attention mechanism, expressive attention (EA), which is based on the presumption that low attention should correspond to orthogonal query and key configurations. This venue allows attention to express itself in the entire attention space, as illustrated in Fig. 1. Interestingly, EA is realized when using $(\mathbf{Q}^{T}\mathbf{K})^{2}$ for calculating attention weights, viz the squared scalar product. For applications, this approach can be implemented with minimal efforts, normally just a single line of code needs to be modified. Computational complexity and memory requirements are not affected.
We present a comparative evaluation of vanilla DPA and EA. For this purpose we employ a suite of autoregressive prediction tasks (NT prediction tasks) that are based on time series generated by the delayed addition of two or more numbers, where addition is modulo a generic basis $N$, like 2 or 16 (binary/hexadecimal case). NT tasks are generalized delayed XOR tasks that become increasingly difficult to solve for larger $N$ and delays $\tau$ (?). Models are expected to achieve not just a good performance, but 100%, which reflects the view that systems can be said to exhibit reasoning errors when achieving less than 100% accuracy for solvable tasks (?). The question is then, how long perfect performance can be maintained when systematically increasing task difficulty, viz complexity.
Refer to caption Refer to caption
Figure 1: Comparing keys (colored arrows) with a given query (black arrow), constitutes the basis of ML attention. Attention weights are large/small when query and key match/mismatch (green/red), or in between, viz neutral (dashed yellow). For the standard dot-product attention (DPA, left panel, see (1)), attention weights are large/small for parallel/antiparallel keys and query. For expressive attention (EA, right panel, defined in (2)), both parallel and antiparallel alignments form matching pairs, with orthogonal keys and queries leading to a mismatch.
2 Related work
Alternatives to the standard, attention-based transformer have been studied intensively during the past years. At a basic level, non-transformer based architectures for sequence processing have been proposed, like structured and/or selective state sequence models (?, ?). Other approaches aim to reduce compute costs, for example by making use of a multi-to-multi recurrent network analogy (?). Another route is linearized attention (?, ?, ?) for which compute scales linearly with context length, and not quadratically. In practice, linear attention comes with various problems of its own, like unbounded gradients and attention dilution (?, ?). In general, linearized attention models substitute the softmax operation of classical dot-product attention by various linear kernel functions, which may have stochastic components (?, ?), well designed spatial dependencies (?), or being derived from a singular value decomposition (?).
A different aspect is addressed by rotary positional embedding (RoPE), which adds positional information directly to the attention mechanism, via position-dependent rotations of query and key vectors (?, ?). RoPE can be used in particular for extending context windows (?). Somewhat similar in spirit is $(IA)^{3}$, which uses element-wise rescaled key and values vectors (in addition to rescaled hidden layer activities) for fine-tuning pretrained LLMs for downstream tasks (?).
Two main routes are available for testing the performance of sequence processing architectures, such as transformers. The first is to use databases relevant for real-world applications (?, ?), the second is to rely on synthetic test suites. The latter approach is used standardly when studying learning biases, e.g., when comparing length generalization scores for functions like ‘parity’, ‘majority’, ‘first’ or ‘mean’ (?, ?). Synthetic test environments have been employed also for the study of attention glitches in reasoning tasks (?). For our studies we use NT tasks, a suite of synthetic autoregressive prediction tasks which can be tuned to a desired level of difficulty, viz complexity. We will study in particular the transition between good and optimal performance when increasing context length.
3 Background
Multiplying token activity with the respective query, key and value matrices generates three vectors, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$ and $\mathbf{V}_{i}$, specifically for each token $i$ in a given attention layer. The activity $\mathbf{y}_{m}$ of token $m$ is given by $\mathbf{y}_{m}=\sum_{k\leq m}a_{mk}\mathbf{V}_{k}$, when masked attention is used. The attention matrix $a_{mk}\geq 0$ is normalized row-wise, $1=\sum_{k}a_{mk}$, encoding how much information is transferred from token $k$ to token $m$. This setup implements information routing for any suitable mechanism determining the individual $a_{mk}$. The standard approach (?),
$a_{mk}\big{|}_{\rm DPA}=\frac{1}{Z_{m}}\exp\left(\beta\mathbf{Q}_{m}^{T}\mathbf{K}_{k}\right)\,$ (1)
takes the scalar product $\mathbf{Q}^{T}\mathbf{K}$ as the fundamental similarity measure. The softmax operation included in (1) transforms the basic similarity measure into a probability distribution function, with $Z_{m}$ being an appropriate normalization factor. One can set $\beta=1$ without affecting performance, as done here, as the standard expression $\beta=1/\sqrt{N_{\rm con}}$ only affects the scaling of entries of the key and query matrices with context length, $N_{\rm con}$ (?). For completeness we tested that the value of $\beta$ does not affect results.
As an alternative to (1) we investigate expressive attention (EA), defined by
$a_{mk}\big{|}_{\rm EA}=\frac{1}{N_{m}}\frac{z_{mk}^{2}}{1+z_{mk}^{2}},\quad\quad z_{mk}=\mathbf{Q}_{m}^{T}\mathbf{K}_{k}\,,$ (2)
where $N_{m}$ is a normalization factor. An illustration is presented in Fig. 1. Importantly, EA is a function of the squared dot-product $z_{mk}$. In leading order, expressive attention is quartic in token activities and hence qualitative different from linear attention, which is bilinear in token activities.
Attention space geometry
On a general level, attention mechanisms based on comparing keys and query vectors do allocate semantic meanings to the respective geometric configurations. Classically, as defined by (1), parallel/anti-parallel keys and queries correspond to high/low attention weights. Attention weights are of intermediate size for orthogonal arrangements.
Full attention matrices tend to be sparse (?), which allows in many instances to reduce the dimensionality of attention matrices from the start via a dedicated decimation process (?, ?). This observation suggest that it may be advantageous if orthogonal configurations would correspond to reduced attention weights. Eq. (2) encodes this concept, as illustrated in Fig. 1. The naming of our approach, ‘expressive attention’, is motivated by above geometric considerations.
Figure 2: For a time series of symbols (boxes), the structure of the NT prediction task (teal), as defined by (4), and of the NT-S variant (red), see (8).
4 Experimental settings
Suite of Tasks
A basic non-linear autoregressive task is to predict the next token of a time series (?). For our investigation we are interested in a suite of prediction tasks which allows to systematically increase and tune difficulty. As a motivation we first consider the basic XOR setting (?).
$x(t)$ $=$ $\mathrm{XOR}\big{(}x(t-1),x(t-2)\big{)}$ (3)
$=$ $\big{[}x(t-1)+x(t-2)\big{]}\%2\,,$
where we used in the second step that the XOR operation correspond to the addition of two boolean inputs $x=0/1$, modulo two. As illustrated in Fig. 2, we generalize (3) to the case of a general delay $\tau\in[1,2,3,..]$ and basis $N\in[2,3,4,..]$,
[…]
which defines the NT prediction task. For the XOR series, recovered for N2T1, two types of cyclic patterns are generated,
$011011011011\dots,\qquad\quad 000000000000\dots\,.$ (5)
There are four initial conditions, 11, 01, 10 and 00, of which the first three give rise to cycles of type 011, with the last leading to the default series. The complexity of the associated prediction task increases systematically when increasing N and/or $\tau$. For example, one has
$512\cdot\underline{120}+64\cdot\underline{60}+8\cdot\underline{30}+1\cdot\underline{15}+1\cdot\underline{1}=65536=16^{4}$ (6)
for N16T3, which states that there are 512/64/8/1/1 cycles of length 120/60/30/15/1. One recovers the $N^{\tau+1}=16^{4}$ possible initial conditions when summing up all cycles together with their respective multiplicities. The most demanding system examined in this study is N16T5, for which $16^{6}=2^{24}\approx 16.8\,\mathrm{M}$ initial conditions exists, together with a corresponding number of distinct cycles.
Transformer details
As a testbed, we use a standard decoder-only transformer architecture with $N_{\rm con}$ context token of embedding dimension $d=N$, where $N$ is the basis of the NT task, see Eq. (4). A straightforward orthogonal token embedding of the $N$ symbols $\{0,1,2,\dots,N\!-\!1\}$ is implemented. Attention is causal. Expressive attention works perfectly with standard positional embedding schemes, which we did however not include. This simplifies the analysis of performance as a function of input length. A single transformer bilayer is used throughout this study, with one attention head per token. Skip connections are present, with layer normalization being performed on entry, separately for the attention and the token-wise, tanh feedforward layer. The width of the feed-forward hidden layer is expanded by a factor four.
Training
Training is performed as a function of epochs, with each epoch consisting of $N_{\rm batch}$ predictions. At the start of each epoch a new random NT series is generated and encoded. The first $N_{\rm con}$ symbols are then loaded into context space. For the underlying sequence we define a batch as the task to predict one-by-one the next $N_{\rm batch}=40$ symbols. A basic SGD optimizer is used during training, with momentum $\mu=0.8$ and a learning rate $\epsilon=0.02$. Our aim is for a testbed that allows to study relative performance, in particular as a function of the complexity of the task, and not to attain optimal performance by fine-tuning meta parameters. If not stated otherwise, results shown are averaged over $16$ independent runs,
Readout & Testing
The readout token is connected via a $dN_{\rm con}\times d$ matrix to the uppermost layer, with the task being to predict the next symbol of a NT time series. For the loss function the basic squared difference is taken, inference is greedy. During training, model performance is evaluated by asking the system to predict $N_{g}=50$ subsequent symbols of $N_{\rm test}=100$ distinct, randomly generated NT series. Once training is finished, larger numbers of $N_{\rm test}$ are used for the final evaluation, at least $N_{\rm test}=10^{4}$. The aim is to achieve 100% accuracy.
Learning strategies
For NT tasks, the symbols to be predicted are determined exclusively by two tokens situated at fixed positions from the right end of the input sequence. As a matter of principle, the complexity of the task would independent of $N$, $\tau$ and $N_{\rm con}$, the context length, if the models tested would focus attention on exactly these two input positions. Our results indicate that this is not the case. Perfect performance is achieved nevertheless regularly, which implies that the respective strategies must be based on exploiting longer-range causal correlations. When using this class of strategies, task complexity does increase exponentially both with the size of the basis and the delay, $N$ and $\tau$, becoming easier on the other hand when increasing context length.
Figure 3: For a NT task with basis $N=16$ and delay $\tau=2$, the performance (lines) and the loss (rescaled by a factor 0.75, circles). Result are for the standard dot-product attention (DPA, maroon), and for expressive attention (EA, blue). Performance is $1/16=0.0625$ (baseline) when predictions are random, and $100\%$ when no errors are made. Shown are results for a context length of 32 token (left panel) and 52/56 token (right panel), with DPA needing 56 token to achieve 100% performance. In this case, EA/DPA shown nearly identical training behavior.
5 Results
A representative result is presented in Fig. 3, where simulations for $N=16$ and $\tau=2$ are shown for both types of attention, EA and DPA. This is a task of moderate complexity, with
$64\cdot\underline{56}+16\cdot\underline{28}+4\cdot\underline{14}+1\cdot\underline{7}+1\cdot\underline{1}=4096=16^{3}\,,$ (7)
compare (6), which means that there are 64/16/4/1/1 periodic patterns of lengths 56/28/14/7/1. Three context lengths are considered, $N_{\rm con}=32/52/56$.
  • 100% performance
    Eventually, both systems achieve 100% accuracy, predicting correctly 100 successive tokens for $10^{4}$ random starting N16T2 sequences.
  • structural traps
    Lowering $N_{\rm con}$, eventually both algorithms become trapped in a structural local minimum. This is particularly evident for $N_{\rm con}=32$, as shown in Fig. 3. Both attention mechanism rapidly improve their performance, reaching a plateau of about $55\%$, which is substantially above baseline. At this point the DPA loss function stops improving, with progress slowing down for EA. The phenomenology seen indicates that the enhanced expressivity of EA allows the system to escape the local trap via a comparatively narrow escape route.
  • apparent emergence
    As a function of system size, the DPA performance rapidly increases from about $55\%$ to $100\%$. This happens between $N_{\rm con}=52$ and $56$, but it may be questionable to interpret this performance jump as an ‘emergence phenomenon’. In fact, what happens is that the structural local trap disappears when enlarging training space successively, an expected behavior.
  • task complexity
    For $\tau=2$, only $409616^{3}=$ 4096 distinct sequences exists for $N=16$. The task is hence of modest complexity, given that during training $N_{\rm batch}=40$ shifted sequences are seen for any single epoch. It is hence not surprising that the problem can be solved within 100 epochs for both models when the context length is large enough. When reducing the number adjustable parameters, both models require however larger training compute, if not failing completely. E.g., for $N_{\rm con}=16$ (not included in Fig. 3), expressive attention still converges to 100% performance, needing however about 2000 epochs.
The results presented in Fig. 3 show in addition that training performance is nearly identical for both attention mechanisms when the problem at hand can be solved easily with available resources.
Figure 4: Results for a NT task with a binary basis, $N=2$, and delay $\tau=5$. The model contains 802 adjustable parameters for a context length $N_{\rm con}=16$, as used. Top: Loss (rescaled by 0.75) and performance. For otherwise identical parameters, dot-product attention (DPA) needs substantially longer to converge than expressive attention (EA). Bottom: Snapshots of the respective attention matrices. Typically, EA (left) covers a larger dynamical range (color-coding in $\log_{10}$ scale), than DPA (right).
Binary sequences
It has been pointed out that an important aspect for our understanding of transformers is the study of their loss landscape (?, ?). In this context we further investigate the structural local minima observed in Fig. 3 by comparing the two attention mechanisms for the smallest possible embedding dimension, $d=N$, which is realized for the binary case $N=2$, compare (4). An additional question here is whether EA is advantageous even in this limit.
We specialize to $\tau=5$, for which there are just two NT sequences, namely the default state ‘00000’, and a cycle of length 63. Together the $63+1=64=2^{6}$ possible initial conditions are covered, which makes N2T5 a seeming trivial task. For $N_{\rm con}=16$, and an embedding dimension $d=N=2$, overall model size is however modest, $N_{\rm model}=802$, which may induce pronounced local minima in the loss landscape. This is indeed the case, as the results presented in Fig. 4 show. Averages over 16 runs have been taken.
For DPA one observes extended training spans during which performance fluctuates around the baseline, 50%, essentially without improving. For EA, the equivalent training stage is substantially shorter. The phenomenology observed indicates that learning is dominated by a stochastic escape process (?). Stochastic escape is present when the additive effect of random fluctuating events allows a system to escape a local loss minimum. However, further studies would be needed to substantiate this hypothesis. Of relevance for our study is the observations that expressive attention may lead to a substantially faster convergence than the classical dot-product mechanism.
Also included in Fig. 4 are sample snapshots of the respective attention matrices. Visual inspection indicates that both models make use of the entire input sequence, viz that attention is not focused predominantly on the two causal token, as defined by (4).
Figure 5: For a NT task with basis $N=16$, delay $\tau=5$, and 128 token, the performance (lines) and the loss (rescaled by a factor 0.75, circles). Qualitatively, the results mirror the ones presented in Fig. 3, both for the standard dot-product attention (DPA, maroon), and for expressive attention (EA, blue). When training for 2000 epochs, models see about 0.5% of the $16.8\cdot 10^{6}$ possible sequences.
Large sequence spaces
The number $N_{\rm seq}=N^{\tau+1}$ of distinct NT sequences has been modest for the results presented in Figs. 3 and 4. As a consequence, exact solutions in the form of learned one-by-one mappings $N_{\rm seq}\to N$ are conceivable, at least as a matter of principle. This is however not any more the case for large $N_{\rm seq}$.
In Fig. 5 we present results for N16T5. Model size is about $4\cdot 10^{5}$ for $N_{\rm con}=128$, in terms of adjustable parameters, which is substantially smaller than the number of distinct NT sequences, $N_{\rm seq}=16^{6}\approx 16.8\cdot 10^{6}$. During training, $N_{\rm train}=0.8\cdot 10^{5}$ training sequences are presented within the first 2000 epochs, which is about 0.5% of $N_{\rm seq}$. It is hence likely that the generative mechanism has been encoded in one form or another once performance is either perfect, or near to 100%. As before, we used $5\cdot 10^{5}$ predictions for the evaluation.
On a qualitative level, the results shown in Figs. 5 and 3 are equivalent. Both attention mechanisms converge rapidly to a heuristic strategy with a non-trivial performance, respectively of the order of 0.55 for N16T2 and 0.2 for N16T5. In both cases, DPA remains stuck in the corresponding local minimum, with EA escaping at a reduced pace.
We did not determine at which context length DPA would converge. For $N_{\rm con}=256$, DPA remains in a local trap, but with an improved performance, of roughly 0.44. Increasing the context length further, to $N_{\rm con}=512$, we find that the training of both DPA and EA becomes unstable when our standard learning parameters are used, namely $\mu=0.8$ and $\epsilon=0.02$. One could stabilize training by adjusting momentum and learning rate, which is however outside the scope of our comparative analysis.
Figure 6: For a mixture of two tasks, N16T2 and N16T2-S, the performance of EA (blue) and DPA (maroon). Compare Fig. 3 for the case when only N16T2 is trained. After 2500 epochs (shaded vertical line), learning speed is reduced by a factor four.
Mixture of tasks
The results presented till now concerned setups containing only a specific single tasks. We performed simulations also for mixture of tasks, finding that expressive attention outperforms classical dot-product attention generically also in this case. We define a new task variant, NT-S,
[…]
which corresponds to summing up the $\tau+1$ preceding token, modulo $N$. The two versions, NT-S and NT are identical for $\tau=1$, compare (4) and Fig. 2, but not for $\tau>1$.
Results for $N_{\rm con}=32$ are given in Fig. 6. During training a 50/50 mixture of N16T2 and N16T2-S autoregressive prediction tasks are presented as such, viz without further information. For a given prompt, which could be either a N16T2 or N16T2-S sequence, the system needs to determine on its own both the delay $\tau$ and the type of the task at hand.
NT-S tasks typically lead to shorter cyclic patterns than the corresponding NT task, which makes them easier to learn. For example, the mean cycle length is 47.6 for N16T2, but only 23.8 for N16T2-S. Both systems, EA and DPA have consequently no problem to achieve 100% accuracy on N16T2-S as a single tasks. When combined with N16T2, both systems achieve still above 99% accuracy for N16T2-S, this time as subtask, as shown in Fig. 6. Also of interest is the reduction of the steady-state N16T2 performance in the mixed-task scenario, from about 0.55 to 0.34, which is also the value achieved by DPA.
In our setup, expressive attention is able to solve both tasks to about 99%, Given that N16T2 is shown only 50% of the time, it is not surprising that learning is prolonged, as evident from Fig. 6. As an experiment we reduced learning speed after 2500 epochs, finding that N16T2 performance increases a bit, by about 0.5%. Interference between competing tasks is reduced at lower learning rates.
Tasks are acquired consecutively in order of difficulty, as shown in Fig. 6, which can be interpreted as an instances of unsupervised curriculum learning (?). Our results can also be seen as an explicit example of successive learning of task quanta, as postulated in the quantization theory of neural scaling (?).
Rare events
As a third variant, we introduce via
[…]
a logical if-statement. The generating algorithm remains deterministic, switching from NT to NT-S when a given condition is fulfilled, here when $x(t\!-\!\tau\!-\!1)=0$. For an observer this corresponds to a rare event that occurs with a probability of 1/16=0.0625 when N=16.
Determining the presence of an isolated hidden logical statement, as defined by (9), can be an exceedingly demanding task. In Fig. 7 results for $N_{\rm con}=64/128$ are presented. For these two context lengths N16T2 is comparatively easy, as shown in Fig. 3, which implies that a certain level of performance should be attainable in any case. For DPA this level about 0.36 for $N_{\rm con}=64$, doubling to 0.72 when the context length is raised to 128.
The data presented in Fig. 7 demonstrate that it is difficult also for expressive attention to isolate a lone if-statement, as defined by (9). Training progress is slow, achieving in the end however an average accuracy of 98.8/99.9%, respectively for $N_{\rm con}=64/128$. As usual, averages over 16 random initial conditions have been taken. The majority of runs achieves however perfect performance. We did not investigate the cause of the non-monotonic events showing up during training, which may be due to periods with an increased clustering of rare events.
Figure 7: A N16T2 task switching conditionally to N16T2-S, as defined by (9). Data as averaged over 16 runs.
6 Discussion
Obtaining encouraging results, we presented a first evaluation of expressive attention (EA), focusing on single-layer models with a single attention head. We also investigate models with two or more layers and/or attention heads, together with different learning speeds and/or momentum. The results presented here are found to be robust. Additional investigation would however be necessary for a full assessment of the potential of EA, e.g. in the realm of natural language processing (?, ?), or within the context of formal languages (?, ?). Generally, we expect that expressive attention will fare at least as well as dot-product attention (DPA), a presumption that is based on the respective design principles. For DPA, large and small attention weights are constrained to a one-dimensional manifold within the space of attention heads, which is not the case for EA. However, the size of the performance boost obtainable when substituting DPA by EA may depend strongly on the use case.
Heuristics
Repeatedly, we observed that initial performance gains flatten out rapidly. This happens in particular for small model sizes, but also for DPA in cases where the available resources are sufficient for EA to solve the problem at hand exactly. We argue that the resulting stationary performance plateau indicates the presence of a local minimum in loss landscape. From this trap models may escape either by embedding the task at hand into into a larger dimensional parameter space, or by improved design. For all cases, the level of the observed stationary performance plateau was independent of initial conditions and training details. This led us to the conclusion that the associated local minimum in loss landscape is structural, viz that it corresponds to a heuristic strategy. To examine how this heuristic strategy works would an interesting research question, which is however beyond the scope of the present study.
Reasoning
Inductive reasoning is one of the big challenges of large language models (?, ?). It is encouraging that one of the building blocks, the logical if-statement included, can be extracted and encoded when effective attention is used. Our analysis, which involved logical task switching within the NT-R framework (9), is however only a first indication and it remains to be seen to which extent EA may raise performance in this field.
Attention space geometry
A somewhat intriguing insight emerging from the present study concerns the nature of the attention mechanism. It is clearly possible that qualitative different realizations of the underlying principle may show somewhat comparative performance levels. What matters in first place is the core working principle, dynamic information routing, and only in second place how this principle is implemented in attention space, e.g. using dot-product or expressive attention.
Acknowledgments
We thank Michael Hahn for discussions.